Students took an Algebra 2/Trig Test with a mean of 87.5% with a standard deviation of 4.5%.
What is the probability that a student scored between 84 and 100%?

Answer by stanbon(75887)

Students took an Algebra 2/Trig Test with a mean of 87.5% with a standard deviation of 4.5%.
What is the probability that a student scored between 84 and 100%?
--------------
z(84) = (84-87.5)/4.5 = -0.7778
---
z(100) = (100-87.5)/4.5 = 2.7778
----
P(84 <= x <=100) = P(-0.7778 <= z <= 2.778) = 0.7789
==============
Cheers,
Stan H.
==============